# ref.bib UUITP-22/24

### v1

@@ 1 Introduction
The standard disk potential is recovered from the small $\epsilon_{i}$ expansion of $W(t,c,\epsilon)$.
In this paper, we build on the relation between toric branes and symplectic cuts to uncover deep connections between equivariant disk potentials and higher dimensional Calabi–Yau geometries. A description of symplectic cuts developed by Braverman [Braverman+1999+85+98] involves a family of CY threefolds parameterized by the open-string moduli, which features the singular manifold (1.1) as a distinguished fibre. For the case of a single toric brane, corresponding to a single cut, Braverman’s construction gives rise to a Calabi–Yau fourfold $X_{4}$, while for two branes it gives rise to a Calabi–Yau fivefold $X_{5}$.
@@
A pair of toric branes in $X_{3}$ is encoded by a pair of symplectic cuts, which is the maximal number allowed for Calabi–Yau threefolds. Braverman’s construction provides a description in terms of a Calabi–Yau fivefold $X_{5}$, presented as a fibration of CY3’s over the parameter space of the two branes. The fivefold geometry encodes the equivariant disk potentials of both branes, in addition to the equivariant quantum cohomology ring of $X_{3}$ itself. We therefore find a hierarchy of CY geometries
@@
In the case of two branes, the divisor $X_{2}$ of $X_{3}$ is replaced by a CY1 $X_{1}$ (arising as the intersection of two such divisors, see Figure 1), whose quantum volume $\mathcal{K}^{D}$ generalizes the role of $\mathcal{H}^{D}$ in (1.2), by encoding $\mathcal{H}^{D}$ for each of the underlying cuts
@@
$\mathcal{H}^{D}(t,c,\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}\,\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}B\Big{(}A_{2}\cdot\epsilon,k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\Big{)}\left(\prod_{j=1}^{k}(-y_{j})^{-\sum_{i}\epsilon_{i}}\right)\\ \times F^{(k)}_{D}\Big{(}A_{2}\cdot\epsilon,\sum_{i}\epsilon_{i},\dots,\sum_{i}\epsilon_{i},k\sum_{i}\epsilon_{i};1+y_{1}^{-1},\dots,1+y_{k}^{-1}\Big{)}$ (1.8)
@@
The relation between symplectic cuts and CY4 makes contact with earlier results in the literature on several fronts. In particular, the fact that disk potentials obey inhomogeneous Picard–Fuchs equations was already observed in the case of the real quintic [Walcher:2006rs, Walcher:2008gho, Morrison:2007bm]. Morever, a relation between open strings and periods of Calabi–Yau fourfolds was observed in [Mayr:2001xk, Lerche:2001cw]. Our results provide a derivation of this correspondence, and generalize it to the equivariant setting.
@@
Finally, we consider the possibility of multiple symplectic cuts of $X_{3}$. Since all hyperplanes must obey the Calabi–Yau condition, and must have nontrivial intersection, the maximum number of simultaneously allowed symplectic cuts is two [^2] We assume that the hyperplanes associated to the cuts are in generic position so that their intersection defines a complex one-dimensional submanifold in $X_{3}$. Notice that one could perform an additional cut along one more hyperplane thus reducing the intersection to a zero-dimensional locus, however, this last cut would necessarily break the Calabi–Yau condition.. A double cut of $X_{3}$ consists of the singular space made up of four top-dimensional strata $X_{3}^{\gtrless,\gtrless}$ glued along common divisors, see Figure 1. The four divisors are further glued along a common divisor, which is a Calabi–Yau onefold $X_{1}$. The quantum volume of $X_{1}$, which we denote as $\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$, is defined by the $\epsilon$-GLSM partition function with target $X_{1}$. Here $c_{1},c_{2}$ parameterize the location of the two cuts, and correspond to open string moduli for two toric Lagrangians in $X_{3}$. We find that the quantum volume $\mathcal{K}^{D}$ admits a universal expression
@@
where $H(x_{1},x_{2},z)$ is a Laurent polynomial in the exponentiated variables $x_{\alpha}=\mathrm{e}^{-c_{\alpha}}$, $z_{a}=\mathrm{e}^{-t^{a}}$ which is closely related to the Hori–Vafa mirror curve of $X_{3}$, and $A_{\alpha}\cdot\epsilon$ and $t\cdot M\cdot\epsilon$ are certain linear combinations of the equivariant parameters $\epsilon_{i}$ (and Kähler moduli $t^{a}$), which we define explicitly in Appendix C.3. It turns out that $\mathcal{K}^{D}$ is the most fundamental among quantum volumes, as it encodes both quantum Lebesgue measures $\mathcal{H}^{D}_{\alpha}$ with $\alpha=1,2$ associated to the two symplectic cuts, as well as the quantum volume of $X_{3}$ itself.
Similarly to the construction of Braverman [Braverman+1999+85+98], the double cut can be realized as a codimension-two submanifold inside a two-parameter family of spaces isomorphic to $X_{3}$. The total space of this fibration is defined as a (CY) fivefold $X_{5}$ together with a projection to $\mathbb{C}^{2}$ such that the generic fiber is $X_{3}$, the fiber over either coordinate axis is isomorphic to one of the two CY4 associated to a single cut, and the fiber over the origin is isomorphic to the singular double cut space. Once again, we observe that equivariant volumes of the top strata of the double cut, ,$\mathcal{F}^{D}_{\gtrless,\gtrless}$ descend from equivariant periods of divisors of the CY5. We use this fact to derive extended Picard–Fuchs equations for the quantum double symplectic cut of $X_{3}$. See the main text for details.
@@ Acknowledgements
It is a pleasure to thank Gaëtan Borot, Andrea Brini and Davide Scazzuso for illuminating discussions. We also extend our gratitude to the organizers of the BPS Dynamics and Quantum Mathematics workshop, held at GGI, where part of this work was conducted. L.C. thanks Johanna Knapp, Joseph McGovern and Emanuel Scheidegger for helpful conversations on the matters of this paper. L.C. also thanks the organizers of The Geometry of Moduli Spaces in String Theory scientific program, hosted by the MATRIX Institute, where this work was completed. The work of L.C. was supported by the ARC Discovery Grant DP210103081. The work of P. L. is supported by the Knut and Alice Wallenberg Foundation grant KAW2020.0307. The research of M. Z. is supported by the VR excellence center grant “Geometry and Physics” 2022-06593.
@@ 2.1 Toric branes, symplectic cuts and equivariance
where $\omega$ is the induced symplectic form on the quotient $X_{3}$ and $H_{\epsilon}$ is the Hamiltonian for the $U(1)^{r+1}$-action. At finite values of $\lambda$, the partition function therefore computes a notion of quantum volume of the Calabi–Yau threefold $X_{3}$.
Remark 2.1.
@@
The main claim of [Cassia:2023uvd] is that quantum symplectic cuts defined by hyperplanes associated to toric $A$-branes compute the disk potential of the corresponding brane. To illustrate this claim, we introduce the ‘equivariant superpotential’
@@
The definition of superpotential in (2.17) contains a shift of the open string modulus, compared to the one given in [Cassia:2023uvd]
@@
This shift is purely a matter of convention, and we make this choice in this paper because $W(t,c,\epsilon)$ written in this way will obey certain differential equations with a suggestive form (analogous equations can be obtained for the equivariant superpotential of [Cassia:2023uvd] upon changing variables). This is the same exact shift that appeared between conventions of [Aganagic:2000gs] and those of [Aganagic:2001nx].
@@
$[W(t,c,\epsilon)]_{0}\propto\text{polynomial in $c$}+\sum_{n>0}w_{n}(\mathrm{e}^{-t},\epsilon)\,\mathrm{e}^{-n\,c}\,.$ (2.20)
@@
up to an overall coefficient which possibly depends on $\epsilon_{i}$. Each term in $W_{\mathrm{disk}}$ represents the contribution from holomorphic worldsheet instantons in a certain relative homology class $\beta\in H_{2}^{\rm rel}(X_{3},L)$ where $L$ is the toric Lagrangian. The set of relevant instanton charges depends on the ‘phase’ of the geometry, i.e. the choice of hyperplanes $(h^{1},h^{2})$, and on the sign of relevant linear combinations of closed and open Kähler moduli $\beta\cdot(t,c)=\sum_{a=1}^{r}\beta_{a}t^{a}+\beta_{r+1}c$.
2.2 Equivariant disk potential of generic toric threefolds
The computation of $A$-brane disk potentials by means of quantum symplectic cuts reviewed above is supported by nontrivial checks in several examples, and by a physical interpretation advanced in [Cassia:2023uvd]. In this section, we will argue that (2.20) is in fact true in general, and therefore quantum symplectic cuts compute disk potentials of toric $A$-branes in arbitrary Calabi–Yau threefolds. On the one hand, this will provide an explanation for all the observations made so far. On the other hand, it will clarify how the equivariant superpotential (2.17) is related to the disk potential computed via open string mirror symmetry [Aganagic:2000gs, Aganagic:2001nx].
Using the integral representation of $\operatorname{\Gamma}$ functions in the integrand of (2.14), the quantum Lebesgue measure can be written in the following form
@@ 2.2 Equivariant disk potential of generic toric threefolds
where $x=\mathrm{e}^{-c}$, $z_{a}=\mathrm{e}^{-t^{a}}$ for $a=1,\dots,r$, and the precise definitions of $A_{\alpha}$ and $M$ depend on the specific details of the geometry, see Appendix C for further details. Here it is important to observe that if we regard $H(x,y,z)=0$ as an equation for the variables $x$ and $y$, the locus of its solutions can be identified with the mirror curve of $X_{3}$ (below we sometimes omit the dependence on the complex moduli $z_{a}$). 555 More precisely (2.22) can be written in terms of the mirror curve for the toric brane, starting from its relation to $\mathcal{K}^{D}$ given in (C.30) and by implementing a suitable change of the integration variable, see Appendix C.5. In the present discussion it is implicit that these redefinitions have been implemented, and we treat $H(x,y)$ as the actual mirror curve. This means that $H(x,y)$ might not coincide with the one that appears in an expression for $\mathcal{K}^{D}$ describing a double cut, in general.
In the integral expression in (2.22) we choose a parametrization of the integrand such that the function $H(x,y,z)$ that appears in the denominator is a monic polynomial in $y$ (with no negative powers of $y$, see Remark 2.3), i.e.
@@
$\mathcal{H}^{D}(t,c,\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}B\Big{(}A_{2}\cdot\epsilon,k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\Big{)}\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\\ \times F^{(k)}_{D}\Big{(}A_{2}\cdot\epsilon,\sum_{i}\epsilon_{i},\dots,\sum_{i}\epsilon_{i},k\sum_{i}\epsilon_{i};1+y_{1}^{-1},\dots,1+y_{k}^{-1}\Big{)}$ (2.24)
where $B(a,b)=\frac{\operatorname{\Gamma}(a)\operatorname{\Gamma}(b)}{\operatorname{\Gamma}(a+b)}$ is Euler’s beta function and $H(x,0,z)=\prod_{j=1}^{k}(-y_{j}(x))$ is the product of all the roots, up to sign. The definition of $F^{(k)}_{D}$ and its most relevant properties are collected in Appendix A. In particular, (2.24) follows directly from (2.22) after applying the integral identity (A.9).
Remark 2.3.
In general, $H(x,y,z)$ as given in (C.17) is not of the form (2.23). To take it into this form one needs to redefine $H$ by factoring out overall constants in $y$ and the lowest power of $y$ or by passing to a universal cover such that one can get rid of all fractional powers of $x,y,z$. This introduces a certain redefinition of $\mathcal{H}^{D}$ which is rather mild. We give a few examples:
  • If $H(x,y,z)$ is polynomial in $y$ but not monic, then one can collect an overall function $f(x,z)y^{r}$ where $r$ is the lowest power of $y$ appearing in the polynomial, i.e. $H(x,y,z)=f(x,z)y^{r}\prod_{j}(y-y_{j}(x,z))$. In this case the normalization can be absorbed in a redefinition of $A_{2}\cdot\epsilon\to A_{2}\cdot\epsilon-r\sum_{i}\epsilon_{i}$ and of the factor outside of the integral in (2.22).  [^6] This term appears as an overall multiplicative factor in the computation of the monodromy. This can be seen by repeating the steps of Appendix B.
@@
Thanks to (2.24), the equivariant superpotential defined in (2.17) can be expressed directly in terms of the roots $y_{j}(x)$ of the mirror curve. Let us introduce the monodromy operator $\Delta_{c}$ defined as
$\Delta_{c}\,f(c):=\frac{1}{2\pi\mathrm{i}}\left(f(c+\pi\mathrm{i})-f(c-\pi\mathrm{i})\right)$ (2.25)
for an arbitrary function $f(c)$. It then follows from the definition of $W(t,c,\epsilon)$ that
In order to prove (2.20) it suffices to compute the terms of order $\epsilon^{0}$ in $\Delta_{c}\mathcal{H}^{D}$ using its expression in terms of Lauricella’s function. Details of this computation can be found in Appendix B.
@@
Let $\alpha_{j}\in\mathbb{Q}$ denote the exponents controlling the asymptotic behavior of the roots $y_{j}(x)$ of the mirror curve equation $H(x,y)=0$ in this phase, namely
@@
In other words, $\alpha_{j}$ are the slopes of external legs in the toric diagram of $X_{3}$. The monodromy of $\mathcal{H}^{D}$ can be computed using the general expression in terms of Lauricella’s function (2.24). Details can be found in Appendix B. We find that the monodromy of the regular part, in the phase $x\to 0$, is

### v2

# ref-v2.bib UUITP-22/24

@@ 1 Introduction
$W(t,c,\epsilon)=\frac{1}{2\pi\mathrm{i}}\int^{c+\pi\mathrm{i}}_{c-\pi\mathrm{i}}\mathcal{H}^{D}(t,c^{\prime},\epsilon)\mathop{}\!\mathrm{d}c^{\prime}\,,$ (1.2)
where $t$ collectively refers to the Kähler moduli of $X_{3}$ while $c$ corresponds to the additional Kähler modulus associated to the cut which leads to $X_{2}$. In the l.h.s. of this equation, $c$ enters the equivariant disk potential as the open-string modulus. The standard non-equivariant disk potential is recovered from the small $\epsilon_{i}$ expansion of $W(t,c,\epsilon)$.
In this paper, we build on the relation between toric branes and symplectic cuts to uncover new connections between equivariant disk potentials and higher dimensional Calabi–Yau geometries. A description of symplectic cuts developed by Braverman [Braverman+1999+85+98] involves a family of CY threefolds parameterized by the open-string moduli, which features the singular manifold (1.1) as a distinguished fiber. For the case of a single toric brane, corresponding to a single cut, Braverman’s construction gives rise to a Calabi–Yau fourfold $X_{4}$, while for two branes it gives rise to a Calabi–Yau fivefold $X_{5}$.
@@
A pair of toric branes in $X_{3}$ is encoded by a pair of symplectic cuts. Braverman’s construction provides a description in terms of a Calabi–Yau fivefold $X_{5}$, presented as a fibration of CY3’s over the parameter space of the two branes. The fivefold geometry encodes the equivariant disk potentials of both branes, in addition to the equivariant quantum cohomology ring of $X_{3}$ itself. We therefore find a hierarchy of CY geometries
@@
where the arrows represent inclusions of the generic fibers. The corresponding GLSM’s also form a hierarchy where the arrows, loosely speaking, correspond to successive symplectic quotients (or gaugings). In the case of two branes, the divisor $X_{2}$ of $X_{3}$ is replaced by a CY1 $X_{1}$ (arising as the intersection of two such divisors, see Figure 1), whose quantum volume $\mathcal{K}^{D}$ generalizes the role of $\mathcal{H}^{D}$ in (1.2), by encoding $\mathcal{H}^{D}$ for each of the underlying cuts
@@
$\mathcal{H}^{D}(t,c,\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}\,\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}\mathrm{B}\Big{(}A_{2}\cdot\epsilon,k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\Big{)}\left(\prod_{j=1}^{k}(-y_{j})^{-\sum_{i}\epsilon_{i}}\right)\\ \times F^{(k)}_{D}\Big{(}A_{2}\cdot\epsilon,\sum_{i}\epsilon_{i},\dots,\sum_{i}\epsilon_{i},k\sum_{i}\epsilon_{i};1+y_{1}^{-1},\dots,1+y_{k}^{-1}\Big{)}$ (1.8)
@@
The relation between symplectic cuts and CY4 makes contact with earlier results in the literature on several fronts. In particular, the fact that disk potentials obey inhomogeneous Picard–Fuchs equations was already observed in the case of the real quintic [Walcher:2006rs, Walcher:2008gho, Morrison:2007bm]. Moreover, a relation between open strings and periods of Calabi–Yau fourfolds was observed in [Mayr:2001xk, Lerche:2001cw]. Our results provide a derivation of this correspondence, and generalize it to the equivariant setting.
@@
Finally, we consider the possibility of performing multiple symplectic cuts on $X_{3}$. We assume that for each cut, the intersection of the two half-spaces, $X_{3}^{<}$ and $X_{3}^{>}$, forms a Calabi–Yau manifold. This corresponds to requiring that the hyperplane associated with the cut contains an affine line parallel to the vector $(1,1,\dots,1)$ in the base of the toric fibration. Under this condition, Braverman’s construction ensures that each resulting fourfold remains Calabi–Yau. We refer to such symplectic cuts as “Calabi–Yau type cuts”, or simply “Calabi–Yau cuts”.
If, furthermore, the hyperplanes are in generic position [^2] This excludes parallel, overlapping hyperplanes, which would lead to quotients by trivially acting $U(1)$ groups and thus to Artin stacks that are not of Deligne–Mumford type. with nontrivial intersection, we deduce that at most two CY symplectic cuts can be performed simultaneously. Moreover, the intersection of all resulting half-spaces defines a complex one-dimensional submanifold in $X_{3}$. [^3] For a general CY $d$-fold $X_{d}$, the maximum number of CY cuts would be $d-1$. An additional cut along one more hyperplane would reduce the intersection to a zero-dimensional locus, but this final cut would necessarily break the Calabi–Yau condition.
A double cut of $X_{3}$ then consists of the singular space made up of four top-dimensional strata $X_{3}^{\gtrless,\gtrless}$ glued along common divisors, see Figure 1. The four divisors are further glued along a common divisor, which is a Calabi–Yau onefold $X_{1}$. The quantum volume of $X_{1}$, which we denote as $\mathcal{K}^{D}(t,c_{1},c_{2},\epsilon)$, is defined by the $\epsilon$-GLSM partition function with target $X_{1}$. Here $c_{1},c_{2}$ parameterize the location of the two cuts, and correspond to open string moduli for two toric Lagrangians in $X_{3}$. We find that the quantum volume $\mathcal{K}^{D}$ admits a universal expression
@@
where $H(x_{1},x_{2},z)$ is a Laurent polynomial in the exponentiated variables $x_{\alpha}=\mathrm{e}^{-c_{\alpha}}$, $z_{a}=\mathrm{e}^{-t^{a}}$ which is closely related to the Hori–Vafa mirror curve of $X_{3}$, and $A_{\alpha}\cdot\epsilon$ and $t\cdot M\cdot\epsilon$ are certain linear combinations of the equivariant parameters $\epsilon_{i}$ (and Kähler moduli $t^{a}$), which we define explicitly in Appendix B.3. It turns out that $\mathcal{K}^{D}$ is the most fundamental among quantum volumes, as it encodes both quantum Lebesgue measures $\mathcal{H}^{D}_{\alpha}$ with $\alpha=1,2$ associated to the two symplectic cuts, as well as the quantum volume of $X_{3}$ itself.
Similarly to the construction of Braverman [Braverman+1999+85+98], the double cut can be realized as a codimension-two submanifold inside a two-parameter family of spaces isomorphic to $X_{3}$. The total space of this fibration is defined as a fivefold $X_{5}$ together with a projection to $\mathbb{C}^{2}$ such that the generic fiber is $X_{3}$, the fiber over either coordinate axis is isomorphic to one of the two CY4 associated to a single cut, and the fiber over the origin is isomorphic to the singular double cut space. The space $X_{5}$ is CY only in the case when both cuts are CY as described in the beginning of this section.
Once again, we observe that equivariant volumes of the top strata of the double cut, ,$\mathcal{F}^{D}_{\gtrless,\gtrless}$ descend from equivariant periods of divisors of the CY5. We use this fact to derive extended Picard–Fuchs equations for the quantum double symplectic cut of $X_{3}$. See the main text for details.
@@ Acknowledgements
It is a pleasure to thank Gaëtan Borot, Andrea Brini and Davide Scazzuso for illuminating discussions. We also extend our gratitude to the organizers of the BPS Dynamics and Quantum Mathematics workshop, held at GGI, where part of this work was conducted. L.C. thanks Johanna Knapp, Joseph McGovern and Emanuel Scheidegger for helpful conversations on the matters of this paper. L.C. also thanks the organizers of The Geometry of Moduli Spaces in String Theory scientific program, hosted by the MATRIX Institute, where this work was completed. We would also like to thank both referees for their careful reading of the manuscript, and especially for the constructive and insightful comments.
The work of L.C. was supported by the ARC Discovery Grant DP210103081. The work of P. L. is supported by the Knut and Alice Wallenberg Foundation grant KAW2020.0307. The research of M. Z. is supported by the VR excellence center grant “Geometry and Physics” 2022-06593.
Data Availability
Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Conflict of Interest
The authors have no conflict of interest to declare.
@@ 2.1 Toric branes, symplectic cuts and equivariance
where $\omega$ is the induced symplectic form on the quotient $X_{3}$ and $H_{\epsilon}$ is the Hamiltonian for the $U(1)^{r+3}$-action. At finite values of $\lambda$, the partition function therefore computes a notion of quantum volume of the Calabi–Yau threefold $X_{3}$.
@@
Solutions of the non-equivariant PF equations are referred to as “periods” due to the fact that by mirror symmetry they can be regarded as periods of the top holomorphic form of the mirror CY. In the equivariant setting, we refer to the solutions of (2.6) as “equivariant periods” even though the precise details of the statement of mirror symmetry are yet to be worked out. An explicit basis of such solutions, is labeled by the torus fixed points in the target $X_{3}$ as discussed in [Cassia:2022lfj], however certain special linear combinations of the basis elements appear naturally by applying difference operators in the Kähler moduli on the quantum volume $\mathcal{F}^{D}$. For each toric divisor in $X_{3}$, we define the difference operator
$\frac{1-\mathrm{e}^{2\pi\mathrm{i}\operatorname{\mathcal{D}}_{i}}}{2\pi\mathrm{i}}\,,$ (2.8)
where the exponential of the differential operator $\operatorname{\mathcal{D}}_{i}$ acts as a finite shift in the variables $t^{a}$,
$\mathrm{e}^{2\pi\mathrm{i}\operatorname{\mathcal{D}}_{i}}f(t^{a})=\mathrm{e}^{2\pi\mathrm{i}\epsilon_{i}}f(t^{a}+2\pi\mathrm{i}Q_{i}^{a})\,.$ (2.9)
We will then define the equivariant period associated to the $i$-th toric divisor as the function
$\Pi(D_{i}):=\frac{1-\mathrm{e}^{2\pi\mathrm{i}\operatorname{\mathcal{D}}_{i}}}{2\pi\mathrm{i}}\cdot\mathcal{F}^{D}(t,\epsilon)$ (2.10)
and similarly, we can define equivariant periods for intersections of multiple divisors by applying multiple difference operators at once. The motivation for such a definition stems from the fact that periods (both equivariant and not) are fully determined by their semiclassical behavior [Cassia:2022lfj], and in the case of the function $\Pi(D_{i})$, it is straightforward to show that the semiclassical behavior is that of the equivariant volume of the divisor $D_{i}$. In particular, one can show that the difference operator (2.8) satisfies two crucial properties. Firstly, it commutes with the PF operators in (2.6), which means that acting with it on a solution automatically gives back another solution. Secondly, it can be expanded as a power series in $\operatorname{\mathcal{D}}_{i}$ with vanishing constant term, which implies that it annihilates all the basis elements associated to fixed points which are not in the divisor $D_{i}$.
@@
The main claim of [Cassia:2023uvd] is that quantum symplectic cuts defined by hyperplanes associated to toric $A$-branes compute the disk potential of the corresponding brane. To illustrate this claim, we introduce the “equivariant disk potential
@@
[…]
$\Delta_{c}\,f(c):=\frac{1}{2\pi\mathrm{i}}\Big{(}f(c+\pi\mathrm{i})-f(c-\pi\mathrm{i})\Big{)}\,,$ (2.21)
it will also be useful in the following to rewrite the defining relation of $W(t,c,\epsilon)$ as the differential-difference equation
@@
The definition of disk potential in (2.20) contains a shift of the open string modulus, compared to the one given in [Cassia:2023uvd]
@@
This shift is purely a matter of convention, and we make this choice in this paper because $W(t,c,\epsilon)$ written in this way will obey certain differential equations with a suggestive form (analogous equations can be obtained for the equivariant disk potential of [Cassia:2023uvd] upon changing variables). This is the same exact shift that appeared between conventions of [Aganagic:2000gs] and those of [Aganagic:2001nx].
@@
$[W(t,c,\epsilon)]_{0}=\text{polynomial in $c$}+\sum_{n>0}w_{n}(\mathrm{e}^{-t},\epsilon)\,\mathrm{e}^{-n\,c}\,.$ (2.25)
@@
up to an overall constant coefficient. Each term in $W_{\mathrm{disk}}$ represents the contribution from holomorphic worldsheet instantons in a certain relative homology class $\beta\in H_{2}^{\rm rel}(X_{3},L)$ where $L$ is the toric Lagrangian. The set of relevant instanton charges depends on the ‘phase’ of the geometry, i.e. the choice of hyperplanes $(h^{1},h^{2})$, and on the sign of relevant linear combinations of closed and open Kähler moduli $\beta\cdot(t,c)=\sum_{a=1}^{r}\beta_{a}t^{a}+\beta_{r+1}c$.
2.2 Equivariant disk potential of general toric threefolds
The computation of $A$-brane disk potentials by means of quantum symplectic cuts reviewed above is supported by nontrivial checks in several examples, and by a physical interpretation advanced in [Cassia:2023uvd]. In this section, we will argue that (2.26) is in fact true in general, and therefore quantum symplectic cuts compute disk potentials of toric $A$-branes in arbitrary Calabi–Yau threefolds. On the one hand, this will provide an explanation for all the observations made so far. On the other hand, it will clarify how the equivariant disk potential (2.20) is related to the disk potential computed via open string mirror symmetry [Aganagic:2000gs, Aganagic:2001nx].
In order to prove a relation similar to (2.26) valid for general toric CY3, we will use (2.22) to argue that, up to overall constants, the instanton part of the non-equivariant expansion of $\partial_{c}W(t,c,\epsilon)$ matches with the derivative of the disk potential $W_{\mathrm{disk}}$. More precisely, we claim that
$\Delta_{c}\left[\mathcal{H}^{D}(t,c,\epsilon)\right]_{0}=\text{polynomial in $c$}+\sum_{j}k_{j}\,\partial_{c}W^{(j)}_{\mathrm{disk}}(\mathrm{e}^{-t},\mathrm{e}^{-c})\,.$ (2.27)
where $[\mathcal{H}^{D}]_{0}$ is defined as the order zero term in $\xi$ in the series expansion of $\mathcal{H}^{D}(t,c,\xi\epsilon)$, analogously to (2.24), while the sum in the r.h.s. ranges over different branches of the mirror curve as explained below. Moreover, $k_{j}$ are certain proportionality constants fixed by the details of the cut. In specific examples, this more general relation reduces to the simpler relation (2.26) observed in [Cassia:2023uvd].
We begin by observing that one can use the integral representation of $\operatorname{\Gamma}$ functions in the integrand of (2.17), to write the quantum Lebesgue measure in the following form
@@ 2.2 Equivariant disk potential of general toric threefolds
where $x=\mathrm{e}^{-c}$, $z_{a}=\mathrm{e}^{-t^{a}}$ for $a=1,\dots,r$, and the precise definitions of $A_{\alpha}$ and $M$ depend on the specific details of the geometry, see (B.43) and its derivation in Appendix B for further details. Here it is important to observe that if we regard $H(x,y,z)=0$ as an equation for the variables $x$ and $y$, the locus of its solutions can be identified with the mirror curve of $X_{3}$ (below we sometimes omit the dependence on the complex moduli $z_{a}$). 666 More precisely (2.28) can be written in terms of the mirror curve for the toric brane, starting from its relation to $\mathcal{K}^{D}$ given in (B.43) and by implementing a suitable change of the integration variable, see Appendix B.5. In the present discussion it is implicit that these redefinitions have been implemented, and we treat $H(x,y)$ as the actual mirror curve. This means that $H(x,y)$ might not coincide with the one that appears in an expression for $\mathcal{K}^{D}$ describing a double cut, in general.
In the integral expression in (2.28) we choose a parametrization of the integrand such that the function $H(x,y,z)$ that appears in the denominator is a monic polynomial in $y$ (with no negative powers of $y$, i.e.
@@
$\mathcal{H}^{D}(t,c,\epsilon)=\operatorname{\Gamma}\Big{(}\sum_{i}\epsilon_{i}\Big{)}\mathrm{B}\Big{(}A_{2}\cdot\epsilon,k\sum_{i}\epsilon_{i}-A_{2}\cdot\epsilon\Big{)}\frac{\mathrm{e}^{-t\cdot M\cdot\epsilon}\,x^{A_{1}\cdot\epsilon}}{H(x,0,z)^{\sum_{i}\epsilon_{i}}}\\ \times F^{(k)}_{D}\Big{(}A_{2}\cdot\epsilon,\sum_{i}\epsilon_{i},\dots,\sum_{i}\epsilon_{i},k\sum_{i}\epsilon_{i};1+y_{1}^{-1},\dots,1+y_{k}^{-1}\Big{)}$ (2.30)
where $\mathrm{B}(a,b)=\frac{\operatorname{\Gamma}(a)\operatorname{\Gamma}(b)}{\operatorname{\Gamma}(a+b)}$ is Euler’s beta function and $H(x,0,z)=\prod_{j=1}^{k}(-y_{j}(x))$ is the product of all the roots, up to sign. The definition of $F^{(k)}_{D}$ and its most relevant properties are collected in Appendix A. In particular, (2.30) follows directly from (2.28) after applying the integral identity (A.9).
In general, $H(x,y,z)$ as given in (B.17) is not of the form (2.29). To bring it into this form one needs to redefine $H$ by factoring out overall constants in $y$ and the lowest power of $y$ or by passing to a universal cover such that one can get rid of all fractional powers of $x,y,z$. This introduces a certain redefinition of $\mathcal{H}^{D}$ which is rather mild. We give a few examples:
  • If $H(x,y,z)$ is polynomial in $y$ but not monic, then one can collect an overall function $f(x,z)y^{r}$ where $r$ is the lowest power of $y$ appearing in the polynomial, i.e. $H(x,y,z)=f(x,z)y^{r}\prod_{j}(y-y_{j}(x,z))$. In this case the normalization can be absorbed in a redefinition of $A_{2}\cdot\epsilon\to A_{2}\cdot\epsilon-r\sum_{i}\epsilon_{i}$ and of the factor outside of the integral in (2.28).  [^7] This term appears as an overall multiplicative factor in the computation of the monodromy. This can be seen by repeating the steps of Appendix C.
@@
Thanks to (2.30), we are able to express the equivariant disk potential defined in (2.20) directly in terms of the roots $y_{j}(x)$ of the mirror curve. Let $\alpha_{j}\in\mathbb{Q}$ denote the exponents controlling the asymptotic behavior of the roots $y_{j}(x)$ of the mirror curve equation $H(x,y)=0$ in this phase, namely
@@
In other words, $\alpha_{j}$ are the slopes of external legs in the toric diagram of $X_{3}$. The monodromy of $\mathcal{H}^{D}$ can be computed using the general expression in terms of Lauricella’s function (2.30), as explained in Appendix C.2. Assuming, for the sake of concreteness, that the roots of the curve satisfy $|1+y_{j}^{-1}|<1$, we can use an explicit series expansion [^8] Power series representations of Lauricella functions are generically not unique, and in fact depend on a choice of region for the arguments of the function within which the series is convergent. See the end of Appendix C.2 for a discussion of this issue. of the Lauricella to compute the monodromy and we find that the regular part, in the phase $x\to 0$, is